How do you set up the problem for (x+4)(x-9)=4x, I know that I need to subtract 4x from each side to set the problem to 0 but I do not know what to do after that. Do I distribute the -4x to (x+4) and then factor out the - and then solve for x?

Answer by stanbon(75887)

set up the problem for
(x+4)(x-9)=4x
-----
x^2-5x-36 = 4x
-----
x^2 - 9x -36 = 0
Factor:
----
(x-12)(x+3) = 0
x = 12 or x = -3
====================
Cheers,
Stan H.